Kacey is picking out jewelry to wear for school. She has 10 rings, 4 bracelets, and 4 necklaces. How many jewelry combinations are possible if she is going to wear 1 ring, 1 bracelets , and 2 necklaces?
Answer by Boreal(15235)

10*4*(4 choose 2)=40*6=240 choices